# Real-Time Learning of Predictive Dynamic Obstacle Models for Robotic Motion Planning

### v1

@@ Abstract
Autonomous systems often must predict the motions of nearby agents from partial and noisy data. This paper asks and answers the question: "can we learn, in real-time, a nonlinear predictive model of another agent’s motions?" Our online framework denoises and forecasts such dynamics using a modified sliding-window Hankel Dynamic Mode Decomposition (Hankel-DMD). Partial noisy measurements are embedded into a Hankel matrix, while an associated Page matrix enables singular-value hard thresholding (SVHT) to estimate the effective rank. A Cadzow projection enforces structured low-rank consistency, yielding a denoised trajectory and local noise variance estimates. From this representation, a time-varying Hankel-DMD lifted linear predictor is constructed for multi-step forecasts. The residual analysis provides variance-tracking signals that can support downstream estimators and risk-aware planning. We validate the approach in simulation under Gaussian and heavy-tailed noise, and experimentally on a dynamic crane testbed. Results show that the method achieves stable variance-aware denoising and short-horizon prediction suitable for integration into real-time control frameworks.
@@ I Introduction
Autonomous robotic systems often operate in environments where other dynamic and non-coordinated agents operate. For example, autonomous cars must navigate around vehicles, pedestrians, and cyclists [1, 2]. In autonomous drone racing, drones must avoid crashing into other drones on the racecourse [3, 4] and in maritime robotics, a shipboard robotic arm or autonomous crane must plan for the payload’s motion while compensating for the ship’s sea-induced oscillations. In such settings, the other agent’s dynamics and intentions are typically unknown while onboard sensing provides noisy, partial observations. Safe, efficient behavior therefore hinges on accurate short-horizon prediction under uncertainty to enable collision-free planning and real-time control. Thus, this paper introduces a data-driven framework for real-time learning and short-horizon prediction from noisy, partial observations.
@@
Different approaches have been developed for integrating prediction and planning in dynamic environments where other agents’ dynamics and intentions are unknown. Classical geometric planners such as the Velocity-Obstacle (VO) family [5], Reciprocal VO (RVO), Optimal Reciprocal Collision Avoidance (ORCA) and Acceleration Velocity Obstacles extensions (AVO)) [6, 7, 8, 9], embed simple relative-motion estimates to predict collisions and select controls outside forbidden sets. These methods are efficient but assume simplified behavior models and near-perfect state estimates. However, real agents often violate these assumptions. For instance, some thrown objects may be roughly ballistic [10], but aerodynamically complex bodies, e.g., frisbees, do not follow a simple ballistic model. Additionally, behavior-driven agents, e.g., pedestrians, require data-driven intent models to capture variability beyond constant-velocity assumptions [11]. Lastly, constraint-driven agents, e.g., autonomous vehicles, have future trajectories restricted by road geometry and traffic rules. Across all these instances, VO-style methods assume perfect knowledge of other agents’ states (exact positions and velocities with instantaneous updates) and neglect measurement noise or sensing delays [12]. This reliance on idealized assumptions highlights the gap between geometric formulations and the stochastic, nonlinear realities of robotics applications.
@@
When other agents’ dynamics are unknown, it is desirable to learn predictive models online from streaming observations. In practice, only partial and noisy measurements are available, while key latent states are unobserved, degrading forecasts and the safety of plans. Prior methods assume full-state observability for simplicity [13]. This includes frameworks that learn dynamics directly from data offline, e.g., classical AutoRegressive Moving Average (ARMA) models [14], modern sequence learners such as Recurrent Neural Networks (RNNs), transformers and tokenized representations [15]. These off-line methods capture complex behavior but require large datasets and retraining [16] and adapt poorly to distribution shift [17]. Online variants update models during deployment, e.g., Kalman filtering, recursive least squares, and online Gaussian processes [18], but typically presume known or structured noise distributions [19].
@@
Koopman theory is a complimentary alternative to learning these dynamic models. Formally, the Koopman operator models a nonlinear dynamical system via a linear operator on a vector space of functions [20]. Since the operator acts on an infinite dimensional vector space, real-time analysis and prediction is intractable. A practical approach is to approximate the operator’s action on a finite-dimensional subspace (which is typically of greater dimension than the state-space of the nonlinear system), enabling the use of well-developed methods from linear algebra and linear system theory. These approximations are often performed via orthogonal projections on the subspace of choice, also referred to as truncations of the operator’s action. Dynamic Mode Decomposition (DMD) [21] and its variant Extended-DMD (EDMD) [22] are well-known examples of such projection-based algorithms. Hankel-DMD uses time-delay embedding to estimate an effective state-space dimension [23, 24]. Such truncated Koopman-based models offer computationally feasible approximations of nonlinear dynamics with proven utility in robotics for prediction and control [25, 26]. However, sensor measurements may be noisy and given that the fidelity of real-time Koopman predictors depends on data quality, noise degrades these learned models.
@@
Existing denoising frameworks to resolve noisy measurements prior to model fitting include classical low-rank denoising techniques i.e., truncated Singular Value Decomposition (tSVD), Principal Component Analysis (PCA) [27] and Proper Orthogonal Decomposition (POD) [28] provide optimal projections but rely on ad hoc rank selection. Structured Hankel denoising [29] enforces temporal consistency but requires tuning and degrades under non-Gaussian noise. Scalable variants like randomized SVD [30] improve efficiency but depend on problem-specific thresholds, while Eigensystem Realization Algorithm (ERA) [31] and Singular Spectrum Analysis (SSA) [32] remain sensitive to noise.
@@
We propose an adaptive, denoising, sliding-window Hankel–DMD framework for real-time prediction from noisy partial measurements. At each step, a finite buffer fits a local model, balancing responsiveness to non-stationary behavior with enough samples for reliability. We first denoise via Cadzow’s structured low-rank projection, with rank $\widehat{r}$ chosen by Singular Value Hard Thresholding (SVHT) on a Page embedding. We prove that, under mild conditions, Page and Hankel have the same finite-sample rank, so the SVHT choice transfers to the Hankel matrix for Cadzow denoising. The resulting denoised Hankel matrices yield a sequence of local models and noise-variance estimates for uncertainty-aware planning. We validate the framework in both simulation and hardware experiments under Gaussian and non-Gaussian correlated heavy-tailed noise and we show that the framework offers robust denoised prediction results suitable for real-time control and planning tasks.
@@
$\mathbb{R}$ and $\mathbb{N}$ denote real and natural numbers.
$I_{n}$ and $0_{m\times n}$ to denote the $n\times n$ identity matrix and $m\times n$ zero matrix, respectively. For matrix $\in\mathbb{R}^{m\times n}, AA^{\top}$, $A^{\dagger}$, and $\operatorname{rank}(A)$ denote its transpose, Moore–Penrose pseudo-inverse, and rank respectively. Given vector $v\in\real^{n}$, we denote by $\operatorname{diag}(v)$, the $n\times n$ diagonal matrix with elements of $v$ on its main diagonal. Given vectors $\{v_{1},\dots,v_{k}\}\subset\real^{n}$, we denote the set comprised of all their linear combinations by $\operatorname{span}\{v_{1},\ldots,v_{k}\}$.
For a random variable $X$, we denote its variance by $\sigma_{X}^{2}$ and we drop the subscript when the context is clear. Given functions $f$ and $g$ with matching domain and co-domain, we denote the composition by $f\circ g(x):=f(g(x))$.
@@ II Problem Definition
Consider a robot that operates in the vicinity of another moving entity $\mathcal{O}$ whose behavior and dynamics are unknown. The robot’s onboard sensors detect $\mathcal{O}$ and provide noisy, partial measurements of the obstacle, denoted by $x_{t}\in\mathbb{R}^{n_{x}}$, which are sampled at a uniform interval $\Delta t\geq 0$, and capture kinematic quantities such as velocities or angular rates. While $x_{t}$ may not directly encode full Cartesian position, it carries sufficient temporal information to enable short-horizon motion prediction. The motion of $\mathcal{O}$ is governed by a discrete-time dynamical system with latent (unobserved) variables and unmeasured control inputs in the form: $z_{t+1}=\hat{f}(z_{t},u_{t})$ where $z_{t}$ denotes the state, which may include latent (unobserved) variables and $u_{t}$ an input from an unknown policy $u_{t}=c(z_{t})$. This induces an autonomous system $f$ as
@@
where $\eta_{t}\in\mathbb{R}^{n_{x}}$ is measurement noise (unknown distribution), and $C$ defines the state-output map. In robotics, this discrete map typically arises from discretization (or discrete sampling) of a continuous-time system. We maintain a sliding buffer of $N$ recent measurements $\mathcal{B}_{t}=\{x_{t-N+1},\ldots,x_{t}\}$ for online processing.
Problem Statement: Given streaming noisy observations $\{x_{t}\}$ from a single trajectory and a sliding data buffer of fixed length $N$, our objective is to construct, in real time, a tractable and adaptive local model. This model should produce multi-step forecasts of future measurements $\{\bar{x}_{t+1},\ldots,\bar{x}_{t+N_{h}}\}$ over a prediction horizon $N_{h}\in\mathbb{N}$, in a manner suitable for downstream tasks such as prediction, planning, and control.
@@ III-A Hankel and Page Matrices
Guided by Takens’ delay-embedding theorem, that the dynamics of a system can be reconstructed from delay-coordinate maps of a single observable when the embedding dimension is sufficiently large [33], we apply delay embedding to the output measurements to extract informative structure from the system. Given a sequence of $N$ measurements $\{x_{i-N+1},\ldots,x_{i}\}\subset\mathbb{R}^{n_{x}}$, we denote the associated (block) Hankel matrix $H^{\,L}_{i-N+1:i}\in\mathbb{R}^{(Ln_{x})\times(N-L+1)}$ by:
$H^{\,L}_{i-N+1:i}\;=\;\begin{bmatrix}x_{i-N+1}&x_{i-N+2}&\cdots&x_{i-L+1}\\ x_{i-N+2}&x_{i-N+3}&\cdots&x_{i-L+2}\\ \vdots&\vdots&\ddots&\vdots\\ x_{i-N+L}&x_{i-N+L+1}&\cdots&x_{i}\end{bmatrix}$ (2)
@@
$P^{\,L}_{i-N+1:i}=\begin{bmatrix}x_{i-N+1}&x_{i-N+L+1}&\cdots&x_{i-L+1}\\ x_{i-N+2}&x_{i-N+L+2}&\cdots&x_{i-L+2}\\ \vdots&\vdots&\ddots&\vdots\\ x_{i-N+L}&x_{i-N+2L}&\cdots&x_{i}\end{bmatrix}$ (3)
@@ III-B Hankel-DMD
Here, we review a variant of Hankel-DMD (cf. [23, 24]) suitable for our problem based on the delay embedding of the observables which measure the output of the dynamical system (II).
Given the measurement sequence $\{x_{0},x_{1},\ldots,x_{n}\}$, form the one step shifted Hankel matrices $H_{0:n-1}^{L}$ and $H_{1:n}^{L}$ according to notation in (2). Note that the columns of $H_{1:n}^{L}$ can be generated by pushing the columns of $H_{0:n-1}^{L}$ one time step forward in time. To estimate the one-step propagator, Hankel-DMD relies on the following least Frobenius-norm problem in the delay coordinates,
@@ III-C Cadzow Algorithm for Denoising
Here, we review a variant of Cadzow’s algorithm for denoising Hankel matrices[34]. In lifted dynamics that are approximately linear (e.g., Koopman-based approximations), noise-free data lives on a low-dimensional vector space. Hence, for sufficiently large delay embedding window $L$, the noise-free Hankel trajectory matrix becomes low rank. However, with measurement noise, the observed Hankel matrix is generally full rank. Formally, let $H^{L}\in\real^{(L\,n_{x})\times(N-L+1)}$ be a Hankel matrix constructed from $N$ consecutive noisy measurements with embedding window $L$. Decompose $H^{L}$ as $H^{L}\;=\;\widehat{H}^{L}\;+\;\Delta$ with
@@
where $\Delta$ denotes measurement noise and $\widehat{H}^{\,L}$ is the noise-free Hankel matrix. The objective is to recover the (unknown) noise-free, low-rank Hankel matrix $\widehat{H}^{\,L}$ from $H^{\,L}$. The Cadzow algorithm tries to achieve this objective by alternating between projecting $H^{L}$ on the set of matrices with rank of at most $r$ and then projecting the result back on the set of Hankel matrices as we explain below.
@@
where $H=U\Sigma V^{\top}$ is the SVD and $\Sigma_{r}$ is created by keeping the top $r$ singular values in $\Sigma$ while setting the rest to zero. The closed-form solution of (7) is a direct consequence of Eckart–Young–Mirsky theorem, see e.g. [35]. Similarly, we define the projection map $\Pi_{\mathsf{H}}$ onto the set of Hankel matrices as
@@
Notably, since the effects of projections $\Pi_{r}$ and $\Pi_{\mathsf{H}}$ can be computed in closed form, the Cadzow algorithm is an attractive choice for real-time denoising, as only a few iterations are sufficient for a reasonable signal-to-noise ratio.

### v2

@@ Abstract
Autonomous systems often must predict the motions of nearby agents from partial and noisy data. This paper asks and answers the question: "can we learn, in real-time, a nonlinear predictive model of another agent’s motions?" Our online framework denoises and forecasts such dynamics using a modified sliding-window Hankel Dynamic Mode Decomposition (Hankel-DMD). Partial noisy measurements are embedded into a Hankel matrix, while an associated Page matrix enables singular-value hard thresholding (SVHT) to denoise the Hankel matrix and estimate its rank. A Cadzow projection enforces structured low-rank consistency, yielding a denoised trajectory and local noise variance estimates. From this representation, a time-varying Hankel-DMD lifted linear predictor is constructed for multi-step forecasts. The residual analysis provides variance-tracking signals that can support downstream estimators and risk-aware planning. We validate the approach in simulation under Gaussian and heavy-tailed noise, and experimentally on a dynamic crane testbed. Results show that the method achieves stable variance-aware denoising and short-horizon prediction suitable for integration into real-time control frameworks.
@@ I Introduction
Autonomous robotic systems often operate in the presence of other dynamic and non-coordinated agents. For example, autonomous cars must navigate around vehicles, pedestrians, and cyclists [18, 51]. In autonomous drone racing, drones must avoid crashing into other drones on the racecourse [42, 21] and in maritime robotics, a shipboard robotic arm or autonomous crane must plan for the payload’s motion while compensating for the ship’s sea-induced oscillations. In such settings, the other agent’s dynamics and intentions are typically unknown, while onboard sensing provides noisy, partial observations of the agent motions. Safe, efficient behavior therefore hinges on accurate short-horizon prediction of agent motion under uncertainty to enable collision-free planning and real-time control. Thus, this paper introduces a data-driven framework for real-time learning and short-horizon prediction from noisy, partial observations.
@@
Different approaches have been developed for integrating prediction and planning in dynamic environments where other agents’ dynamics and intentions are unknown. Classical geometric planners such as the Velocity-Obstacle (VO) family [12], Reciprocal VO (RVO), Optimal Reciprocal Collision Avoidance (ORCA) and Acceleration Velocity Obstacles extensions (AVO)) [46, 47, 19, 48], embed simple relative-motion estimates to predict collisions and select controls outside forbidden sets. These methods are efficient but assume simplified behavior models and near-perfect state estimates. However, real agents often violate these assumptions. For instance, some thrown objects may be roughly ballistic [16], but aerodynamically complex bodies, e.g., frisbees, do not follow a simple ballistic model. Additionally, behavior-driven agents, e.g., pedestrians, require data-driven intent models to capture variability beyond constant-velocity assumptions [6]. Lastly, the future trajectories of constraint-driven agents, e.g., autonomous vehicles, may be restricted by road geometry and traffic rules. Across all these examples, VO-style methods assume perfect knowledge of the moving agents’ states and neglect measurement noise or sensing delays [24]. This reliance on idealized assumptions highlights the gap between geometric formulations and the stochastic, nonlinear realities of robotics applications.
@@
When other agents’ dynamics are unknown, it is desirable to learn predictive models online from streaming observations. In practice, only partial and noisy measurements are available, while key latent states are unobserved, degrading forecasts and the safety of plans. Prior methods assume full-state observability for simplicity [45]. This includes frameworks that learn dynamics directly from data offline, e.g., classical AutoRegressive Moving Average (ARMA) models [26], modern sequence learners such as Recurrent Neural Networks (RNNs), transformers, and tokenized representations [44]. These off-line methods capture complex behavior but require large datasets and retraining [39] and adapt poorly to distribution shift [38]. Online variants (e.g., Kalman Filtering (KF), recursive least squares, and online Gaussian processes [31]) update models during deployment, but presume known or structured noise distributions [37].
@@
Koopman theory is a complimentary alternative to learning these dynamic models. Formally, the Koopman operator models a nonlinear dynamical system via a linear operator on a vector space of functions [33]. Since the operator acts on an infinite dimensional vector space, real-time analysis and prediction is intractable. A practical approach is to approximate the operator’s action on a finite-dimensional subspace (which is typically of greater dimension than the state-space of the nonlinear system), enabling the use of well-developed methods from linear algebra and linear system theory. These approximations are often performed via orthogonal projections on the subspace of choice, also referred to as truncations of the operator’s action. Dynamic Mode Decomposition (DMD) [40] and its variant Extended-DMD (EDMD) [50] are well-known examples of such projection-based algorithms. Hankel-DMD uses time-delay embedding to estimate an effective state-space dimension [3, 29]. Such truncated Koopman-based models offer computationally feasible approximations of nonlinear dynamics with proven utility in robotics for prediction and control [34, 13]. However, sensor measurements may be noisy, and since the fidelity of real-time Koopman predictors depends on data quality, noise degrades these learned models.
@@
Existing denoising frameworks to resolve noisy measurements prior to model fitting include classical low-rank denoising techniques i.e., truncated Singular Value Decomposition (tSVD), Principal Component Analysis (PCA) [1] and Proper Orthogonal Decomposition (POD) [4]. They provide optimal projections but rely on ad hoc rank selection. Structured Hankel denoising [52] enforces temporal consistency but requires tuning and degrades under non-Gaussian noise. Scalable variants like randomized SVD [20] improve efficiency but depend on problem-specific thresholds, while Eigensystem Realization Algorithm (ERA) [27] and Singular Spectrum Analysis (SSA) [17] remain sensitive to noise.
@@
We propose an adaptive, denoising, sliding-window Hankel–DMD framework for real-time prediction from noisy partial measurements. At each time step, a finite data buffer fits a local model, balancing responsiveness to non-stationary behavior with enough samples for reliability. We first denoise via Cadzow’s low-rank projection, with rank $\widehat{r}$ chosen by Singular Value Hard Thresholding (SVHT) on a Page matrix embedding. We prove that, under mild conditions, Page and Hankel matrices have the same finite-sample rank, so the SVHT rank transfers to the Hankel matrix for Cadzow denoising. The resulting denoised Hankel matrices yield a sequence of local models and noise-variance estimates for uncertainty-aware planning. We validate the framework in both simulation and hardware experiments under Gaussian and non-Gaussian correlated heavy-tailed noise. We show that the framework offers robust denoised predictions suitable for real-time control and planning tasks.
@@
$\mathbb{R}$ and $\mathbb{N}$ denote real and natural numbers. $I_{n}$ and $0_{m\times n}$ denote the $n\times n$ identity and $m\times n$ zero matrices, respectively. For matrix $\in\mathbb{R}^{m\times n}, AA^{\top}$, $A^{\dagger}$, and $\operatorname{rank}(A)$ denote its transpose, Moore–Penrose pseudo-inverse, and rank respectively. We denote by $\operatorname{diag}(v)$, the $n\times n$ diagonal matrix with elements of $v\in\mathbb{R}^{n}$ on its main diagonal. All linear combinations of vectors $\{v_{1},\dots,v_{k}\}\subset\mathbb{R}^{n}$ is denoted by $\operatorname{span}\{v_{1},\ldots,v_{k}\}$. We denote the variance of random variable $X$ by $\sigma_{X}^{2}$ and drop the subscript when the context is clear. For functions $f$ and $g$ with matching domain and co-domain, $f\circ g(x):=f(g(x))$ denotes their composition.
@@ II Problem Definition
Consider a robot that operates in the vicinity of another moving entity $\mathcal{O}$ whose behavior and dynamics are unknown. The robot’s onboard sensors detect $\mathcal{O}$ and provide noisy, partial measurements of the obstacle state, denoted by $x_{t}\in\mathbb{R}^{n_{x}}$, which are sampled at a uniform interval $\Delta t\geq 0$, and capture kinematic quantities such as velocities or angular rates. While $x_{t}$ may not directly encode full Cartesian position, it carries sufficient temporal information to enable short-horizon motion prediction. The motion of $\mathcal{O}$ is governed by a discrete-time dynamical system with latent (unobserved) variables and unmeasured control inputs in the form: $z_{t+1}=\hat{f}(z_{t},u_{t})$ where $z_{t}$ denotes the state, which may include latent (unobserved) variables and $u_{t}$ an input from an unknown policy $u_{t}=c(z_{t})$. Under these assumptions, we have an autonomous system
@@
where measurement noise $\eta_{t}\in\mathbb{R}^{n_{x}}$ has an unknown distribution, and $C$ is the state-output map. This discrete time model generally arises from discretization of a continuous-time system. We maintain a sliding buffer of $N$ recent measurements $\mathcal{B}_{t}=\{x_{t-N+1},\ldots,x_{t}\}$ for online processing.
Problem Statement: Given streaming noisy observations $\{x_{t}\}$ from a single trajectory and a sliding data buffer of fixed length $N$, construct, in real time, a tractable and adaptive local model. This model generates multi-step forecasts $\{\bar{x}_{t+1},\ldots,\bar{x}_{t+N_{h}}\}$ over a prediction horizon $N_{h}\in\mathbb{N}$, suitable for downstream planning and control.
@@ III-A Hankel and Page Matrices
Guided by Takens’ delay-embedding theorem (the dynamics of a system can be reconstructed from delay-coordinate maps of a single observable when the embedding dimension is sufficiently large [43]), we apply delay embedding to the obstacle measurements to extract informative information about its dynamics. Given a sequence of $N$ measurements $\{x_{i-N+1},\ldots,x_{i}\}\subset\mathbb{R}^{n_{x}}$, we denote the associated (block) Hankel matrix $H^{\,L}_{i-N+1:i}\in\mathbb{R}^{(Ln_{x})\times(N-L+1)}$ by:
$H^{\,L}_{i-N+1:i}\;=\;\begin{bmatrix}x_{i-N+1}&x_{i-N+2}&\cdots&x_{i-L+1}\\ x_{i-N+2}&x_{i-N+3}&\cdots&x_{i-L+2}\\ \vdots&\vdots&\ddots&\vdots\\ x_{i-N+L}&x_{i-N+L+1}&\cdots&x_{i}\end{bmatrix}$. (2)
@@
$P^{\,L}_{i-N+1:i}=\begin{bmatrix}x_{i-N+1}&x_{i-N+L+1}&\cdots&x_{i-L+1}\\ x_{i-N+2}&x_{i-N+L+2}&\cdots&x_{i-L+2}\\ \vdots&\vdots&\ddots&\vdots\\ x_{i-N+L}&x_{i-N+2L}&\cdots&x_{i}\end{bmatrix}$. (3)
@@ III-B Hankel-DMD
We review a variant of Hankel-DMD (cf. [3, 29]) suitable for our problem. It uses delay embedding on the measurements of obstacle’s observables  (II). Given a measurement sequence $\{x_{0},x_{1},\ldots,x_{n}\}$, form the one step shifted Hankel matrices $H_{0:n-1}^{L}$ and $H_{1:n}^{L}$ according to the Eq.  (2). Note that the columns of $H_{1:n}^{L}$ are shifted one time step forward from the columns of $H_{0:n-1}^{L}$. To estimate the one-step dynamic prediction propagator, $A^{*}$, Hankel-DMD relies on the following least Frobenius-norm problem,
@@ III-C Cadzow Algorithm for Denoising
Here, we review a variant of Cadzow’s algorithm for denoising Hankel matrices[5]. In the approximately linear lifted dynamics (e.g., Koopman-based approximations), noise-free data resides on a low-dimensional vector space. Hence, for a sufficiently large delay embedding window $L$, the noise-free Hankel trajectory matrix becomes low rank. However, with measurement noise, the observed Hankel matrix is generally full rank. Let Hankel matrix $H^{L}\in\real^{(L\,n_{x})\times(N-L+1)}$ be constructed from $N$ consecutive noisy measurements with embedding window $L$. Decompose $H^{L}$ as $H^{L}\;=\;\widehat{H}^{L}\;+\;\Delta$ with
@@
where $\Delta$ denotes measurement noise and $\widehat{H}^{\,L}$ is the noise-free Hankel matrix. The goal is to recover the (unknown) noise-free, low-rank Hankel matrix $\widehat{H}^{\,L}$ from $H^{\,L}$. Cadzow’s algorithm aims for this objective by alternating projections of $H^{L}$ on the set of matrices with rank of at most $r$ with projections of the result back onto the set of Hankel matrices.
@@
where $H=U\Sigma V^{\top}$ is the SVD and $\Sigma_{r}$ is created by keeping the top $r$ singular values in $\Sigma$ while setting the rest to zero. The closed-form solution of (7) is a direct consequence of Eckart–Young–Mirsky theorem  [11]. Similarly, we define the projection map $\Pi_{\mathsf{H}}$ onto the set of Hankel matrices as
@@
Notably, since projections $\Pi_{r}$ and $\Pi_{\mathsf{H}}$ can be computed in closed form, the Cadzow algorithm is an attractive choice for real-time denoising, as only a few iterations are sufficient to yield a reasonable signal-to-noise ratio.